Calculate the acceleration of earth to meet a 1500-kg object. The mass of the Earth is 5.97 X 10²⁴ kg and the radius of Earth is 6.38 x 10⁶ m.


1
Expert's answer
2021-03-29T09:01:54-0400

"F=G\frac{Mm}{r^2}=6.67\cdot10^{-11}\cdot\frac{5.97\cdot10^{24}\cdot1500}{(6.38\cdot10^6)^2}=14674\ (N)"


"a_{Earth}=F/M=\frac{14674}{5.97\cdot10^{24}}=2.5\cdot10^{-21} \ (m/s^2)" . Answer